OpenStudy (anonymous):

A rectangular tank measure 4m long,2 m wide, and 4.8 m high.initially it is half filled with water.Find the depth of water in the tank after 4000 liters more of water are added to it.

OpenStudy (anonymous):

@thomas5267 @s2 @hartnn @harishdeen @arabpride @rane @CO_oLBoY @CGGURUMANJUNATH @chakaron13 @Miracrown help me

OpenStudy (anonymous):

@Kainui @Miracrown

OpenStudy (anonymous):

Change 4000 liters to cu/meters: 4000/1000 = 4 cu/m Find the initial amt of water in the tank, half full would be: (half the given height) 4 * 2 * 2.4 = 19.2 cu/meters Total after 4 cu/m is added to it 19.2 + 4 = 23.2 cu/m total Find the depth (d) 4 * 2 * d = 23.2 8d = 23.2 d = 23.2%2F8 d = 2.9 meter is the depth Check solution: 4 * 2 * 2.9 = 23.2